3. (6) Ann uses the cord to draw a circle of radius 4.80 m ± 0.05 m in a field. Calculate the area of the circle drawn, giving the answer as A ± ΔΑ (area with uncertainty).
Answer #1

The area is given by
   A=\pi r^2
With the uncertainty
  \ln A=\ln \pi +2\ln r
\Rightarrow \frac{dA}{A}=2\frac{dr}{r}
  \Rightarrow {\Delta A}=2A\frac{\Delta r}{r}
And given the values
   r=4.80~m,~~\Delta r = 0.05~m
We calculate the area as
A=\pi r^2=\pi\times (4.8)^2=72.38~m^2
And the uncertainty in the area is
\Rightarrow {\Delta A}=2\times 72.38\times \frac{0.05}{4.8}
   \Rightarrow {\Delta A}=1.51~m^2
So, the final result of the area is
   (72.38\pm 1.51)~m^2